Solve the following using matrix algebra

2X+Y-Z=5

3X-2Y+2Z=-3

X-3Y-3Z=-2


1
Expert's answer
2022-06-13T15:56:39-0400
"A=\begin{pmatrix}
 2 & 1 & -1 \\
 3 & -2 & 2 \\
 1 & -3 & -3 \\
\end{pmatrix}, X=\begin{pmatrix}
 x \\
 y \\
 z \\
\end{pmatrix}, B=\begin{pmatrix}
 5 \\
 -3 \\
 -2 \\
\end{pmatrix}"

"AX=B"

"A^{-1}AX=A^{-1}B"

"X=A^{-1}B"

Augment the matrix with the identity matrix


"\begin{pmatrix}
 2 & 1 & -1& & 1 & 0 & 0 \\
 3 & -2 & 2 & & 0 & 1 & 0 \\
 1 & -3 & -3 & & 0 & 0 & 1 \\
\end{pmatrix}"

"R_1=R_1\/2"


"\begin{pmatrix}
 1 & 1/2 & -1/2& & 1/2 & 0 & 0 \\
 3 & -2 & 2 & & 0 & 1 & 0 \\
 1 & -3 & -3 & & 0 & 0 & 1 \\
\end{pmatrix}"

"R_2=R_2-3R_1"


"\begin{pmatrix}
 1 & 1/2 & -1/2& & 1/2 & 0 & 0 \\
 0 & -7/2 & 7/2 & & -3/2 & 1 & 0 \\
 1 & -3 & -3 & & 0 & 0 & 1 \\
\end{pmatrix}"

"R_3=R_3-R_1"


"\begin{pmatrix}
 1 & 1/2 & -1/2& & 1/2 & 0 & 0 \\
 0 & -7/2 & 7/2 & & -3/2 & 1 & 0 \\
 0 & -7/2 & -5/2 & & -1/2 & 0 & 1 \\
\end{pmatrix}"

"R_2=-2R_2\/7"


"\begin{pmatrix}
 1 & 1/2 & -1/2& & 1/2 & 0 & 0 \\
 0 & 1 & -1 & & 3/7 & -2/7 & 0 \\
 0 & -7/2 & -5/2 & & -1/2 & 0 & 1 \\
\end{pmatrix}"

"R_1=R_1-R_2\/2"


"\begin{pmatrix}
 1 & 0& 0 & & 2/7 & 1/7 & 0 \\
 0 & 1 & -1 & & 3/7 & -2/7 & 0 \\
 0 & -7/2 & -5/2 & & -1/2 & 0 & 1 \\
\end{pmatrix}"

"R_3=R_3+7R_2\/2"


"\begin{pmatrix}
 1 & 0& 0 & & 2/7 & 1/7 & 0 \\
 0 & 1 & -1 & & 3/7 & -2/7 & 0 \\
 0 & 0 & -6 & & 1 & -1 & 1 \\
\end{pmatrix}"

"R_3=-R_3\/6"


"\begin{pmatrix}
 1 & 0& 0 & & 2/7 & 1/7 & 0 \\
 0 & 1 & -1 & & 3/7 & -2/7 & 0 \\
 0 & 0 & 1 & & -1/6 & 1/6 & -1/6 \\
\end{pmatrix}"

"R_2=R_2+R_3"


"\begin{pmatrix}
 1 & 0& 0 & & 2/7 & 1/7 & 0 \\
 0 & 1 & 0 & & 11/42 & -5/42 & -1/6 \\
 0 & 0 & 1 & & -1/6 & 1/6 & -1/6 \\
\end{pmatrix}"

We are done. On the left is the identity matrix. On the right is the inverse matrix.


"A^{-1}=\begin{pmatrix}
 2/7 & 1/7 & 0 \\
11/42 & -5/42 & -1/6 \\
 -1/6 & 1/6 & -1/6 \\
\end{pmatrix}"

"X=\begin{pmatrix}
 x \\
 y \\
 z \\
\end{pmatrix}=\begin{pmatrix}
 2/7 & 1/7 & 0 \\
11/42 & -5/42 & -1/6 \\
 -1/6 & 1/6 & -1/6 \\
\end{pmatrix}\begin{pmatrix}
 5\\
 -3 \\
-2 \\
\end{pmatrix}"

"=\begin{pmatrix}
 10/7-3/7\\
55/42+15/42+14/42 \\
-5/6-3/6+2/6 \\
\end{pmatrix}=\begin{pmatrix}
1\\
2\\
-1 \\
\end{pmatrix}"

"(1, 2, -1)"